Sin^2(theta) = 1/2
Answer by lwsshak3(11628)

Sin^2(theta) = 1/2
For (0,π)
sin^2x=1/2
sinx=1/√2=√2/2
x=π/4, 3π/4 (in quadrants I and II where sin>0)